# Modified scattering for the three dimensional Maxwell-Dirac system

### v4

@@ 4.5. Bounds for a mix of derivatives and vector fields
    (4.24) $\|\Omega\partial^{J}\phi\|_{L^{p}(C_{T}^{+})}\lesssim\|\Omega^{\leq 3}\phi\|_{L^{2}(C_{T}^{+})}^{\frac{1}{3}}\|\partial^{\leq\frac{3}{2}|J|}\phi\|_{L^{q}(C_{T}^{+})}^{\frac{2}{3}}.$
@@
    (4.25) $\|\Omega^{2}\partial^{J}\phi\|_{L^{p}(C_{T}^{+})}\lesssim\|\Omega^{\leq 3}\phi\|_{L^{2}(C_{T}^{+})}^{\frac{2}{3}}\|\partial^{\leq 3|J|}\phi\|_{L^{q}(C_{T}^{+})}^{\frac{1}{3}}.$
@@
(4.27) $T_{1}^{\frac{1}{2}}\|F_{1,I}\|_{L^{2}(C_{T_{1}})}\lesssim\varepsilon\,(\|\Gamma^{\leq 9}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}+\|\Gamma^{\leq 9}A\|_{L^{\infty}L^{3}(C_{T_{1}})}),$
(4.28) $T_{1}^{\frac{1}{3}}\|G_{1,I}\|_{L^{1}L^{\frac{3}{2}}(C_{T_{1}})}\lesssim\varepsilon\,\|\Gamma^{\leq 9}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}.$
@@
$T_{1}^{\frac{1}{2}}\|\partial^{I-I_{1}}A_{\mu}\gamma^{\mu}\Omega\partial^{I_{1}}\psi\|_{L^{2}(C_{T_{1}})}\lesssim T_{1}^{\frac{1}{2}}\|\partial^{I-I_{1}}A\|_{L^{p_{1}}(C_{T_{1}})}\|\Omega\partial^{I_{1}}\psi\|_{L^{q_{1}}(C_{T_{1}})}.$
@@
$\|\partial^{I-I_{1}}A\|_{L^{p_{1}}(C_{T_{1}})}$ $\lesssim\|\partial^{\leq\frac{3}{2}|I|}A\|_{L^{3}(C_{T_{1}})}^{\theta_{1}}\|A\|_{L^{\infty}(C_{T_{1}})}^{1-\theta_{1}}$
$\lesssim T_{1}^{\frac{1}{3}\theta_{1}}\|\partial^{\leq 9}A\|_{L^{\infty}L^{3}(C_{T_{1}})}^{\theta_{1}}(C_{0}\varepsilon{T_{1}}^{-1})^{1-\theta_{1}},$
@@
$\|\Omega\partial^{I_{1}}\psi\|_{L^{q_{1}}(C_{T_{1}})}\lesssim\|\Omega^{\leq 3}\psi\|_{L^{2}(C_{T_{1}})}^{\frac{1}{3}}\|\partial^{\leq\frac{3}{2}|I_{1}|}\psi\|_{L^{q_{2}}(C_{T_{1}})}^{\frac{2}{3}}.$
@@
$\|\partial^{\leq\frac{3}{2}|I_{1}|}\psi\|_{L^{q_{2}}(C_{T_{1}})}\lesssim\|\partial^{\leq\frac{3}{2}|I|}\psi\|_{L^{2}(C_{T_{1}})}^{\theta_{2}}\|\psi\|_{L^{6}(C_{T_{1}})}^{1-\theta_{2}}.$
@@
$\|\Omega\partial^{I_{1}}\psi\|_{L^{q_{1}}(C_{T_{1}})}$ $\lesssim T_{1}^{\frac{1}{6}}\|\Omega^{\leq 3}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}^{\frac{1}{3}}T_{1}^{\frac{1}{2}\theta_{2}\cdot\frac{2}{3}}\|\partial^{\leq 9}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}^{\frac{2}{3}\theta_{2}}$
@@
(4.29) $T_{1}^{\frac{1}{2}}\|\partial^{I-I_{1}}A_{\mu}\gamma^{\mu}\Omega\partial^{I_{1}}\psi\|_{L^{2}(C_{T_{1}})}\lesssim C_{0}\varepsilon(\|\Gamma^{\leq 9}\psi\|_{L^{\infty}L^{2}(C_{T_{1}})}+\|\Gamma^{\leq 9}A\|_{L^{\infty}L^{3}(C_{T_{1}})}).$
@@
$T_{1}^{\frac{1}{2}}\|\partial^{I-I_{1}}A_{\mu}\gamma^{\mu}\partial_{\nu}\Omega\partial^{I_{2}}\psi\|_{L^{2}(C_{T_{1}})}\lesssim T_{1}^{\frac{1}{2}}\|\partial^{I-I_{1}}A\|_{L^{p_{1}}(C_{T_{1}})}\|\partial_{\nu}\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}(C_{T_{1}})}$
@@
$\|\partial_{\nu}\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}(C_{T_{1}})}\lesssim\|\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}(C_{T_{1}})}^{\frac{1}{2}}\|\partial^{\leq 2}\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}(C_{T_{1}})}^{\frac{1}{2}}.$
@@
$\|\partial_{\nu}\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}(C_{T_{1}})}\lesssim\|\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}(C_{T_{1}})}+\|\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}(C_{T_{1}})}^{\frac{1}{2}}\|\partial^{2}\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}(C_{T_{1}})}^{\frac{1}{2}}.$
Using again that the commutator of $\Omega$ with regular derivatives yields regular derivatives which can be dealt with by Gagliardo-Nirenberg estimates, it remains to estimate $\|\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}(C_{T_{1}})}$ and $\|\Omega\partial^{2}\partial^{I_{2}}\psi\|_{L^{q_{1}}(C_{T_{1}})}$. But as $|I_{2}|$ is even, we can proceed here as in the case $|I_{1}|$ even. In conclusion, we also obtain (4.29) in the case $|I_{1}|$ odd.
@@ 5. Pointwise bounds
Our main result here is linear and applies at a fixed dyadic scale $T$. Because of this, we omit the $T^{C\varepsilon}$ factor in Proposition 4.1. Then we want to show that
@@
(5.2) $|\psi|\lesssim t^{-\frac{3}{2}}\langle t-r\rangle_{-}^{-\delta},$
@@
Here $\mathscr{T}$ stands for normalized derivatives in directions which are tangent to the light cones $\{t-r=const\}$. These are spanned by $\mathscr{T}=\{\partial_{r}+\partial_{t},/\!\!\!\!\nabla\}$ where $\!\!\nabla$ stands for the derivatives in angular directions. Equivalently, in the above theorem one may use derivatives tangent to the hyperboloids $\{t^{2}-r^{2}=const\}$.
@@
(5.8) $T^{\frac{1}{2}}\|\Gamma^{\leq 2k}F\|_{L^{2}(C_{T})}+T^{\frac{1}{3}}\|\Gamma^{\leq 2k}G\|_{L^{\frac{3}{2}}(C_{T})}\lesssim 1.$
@@
(5.15) $\|\Gamma^{\leq k}\psi_{+}\|_{L^{2}(C_{TS})}\lesssim$ $\ T^{\frac{1}{2}}\|\Gamma^{\leq k}\psi\|_{X_{T}},$
$\|\Gamma^{\leq k}\psi_{-}\|_{L^{2}(C_{TS})}\lesssim$ $\ S^{\frac{1}{2}}\|\Gamma^{\leq k}\psi\|_{X_{T}}.$
@@
(5.16) $\|P^{\theta}_{+}\Gamma^{\leq k}\psi\|_{L^{2}(C_{TS})}\lesssim T^{\frac{1}{2}}\|{\Gamma^{\leq k}}\psi\|_{X_{T}},$
@@
(5.17) $\|P_{-}^{\theta}\Gamma^{\leq k}\psi\|_{L^{2}(C_{TS})}\lesssim S^{\frac{1}{2}}\|\Gamma^{\leq k}\psi\|_{X_{T}}.$
@@
(5.21) $\|\Gamma^{\leq 6}F_{1}\|_{L^{2}(C_{TS})}\lesssim T^{-\frac{1}{2}}.$
@@
(5.23) $\|\Gamma^{\leq 9}\psi_{-}\|_{L^{2}(C_{TS})}\lesssim S^{\frac{1}{2}},\quad\|\Gamma^{\leq 9}\psi_{+}\|_{L^{2}L^{2}(C_{TS})}\lesssim T^{\frac{1}{2}},\quad\|\Gamma^{\leq 6}F_{\pm}\|_{L^{2}L^{2}(C_{TS})}\lesssim T^{-\frac{1}{2}}.$
@@
(5.24) $\|\Gamma^{\leq 6}\psi_{\pm}\|_{L^{2}(C_{TS})}\lesssim ST^{-\frac{1}{2}}.$
A direct application of (5.22) and (5.23) yields the required estimate for $\Gamma^{\leq 6}\psi_{-}$, but only $\|\Gamma^{\leq 6}\psi_{+}\|_{L^{2}(C_{TS})}\lesssim S^{\frac{1}{2}}$. To improve on that bound, we look at the first equation in (5.22) for $\Gamma^{\leq 7}\psi_{-}$ and note that the corresponding right-hand side is given by
@@
$\|r^{-1}\Omega\Gamma^{\leq 7}\psi_{-}\|_{L^{2}(C_{TS})}$ $\lesssim\frac{1}{T}\|\partial\Omega\Gamma^{\leq 6}\psi_{-}\|_{L^{2}(C_{TS})}+T^{-\frac{1}{2}}$
$\lesssim\frac{1}{T}\|\partial^{\leq 2}\Omega\Gamma^{\leq 6}\psi_{-}\|_{L^{2}(C_{TS})}^{\frac{1}{2}}\|\Omega\Gamma^{\leq 6}\psi_{-}\|_{L^{2}(C_{TS})}^{\frac{1}{2}}+T^{-\frac{1}{2}}$
$\lesssim S^{\frac{1}{4}}T^{-\frac{1}{2}}\|T^{-1}\partial^{\leq 2}\Omega\Gamma^{\leq 6}\psi_{-}\|_{L^{2}(C_{TS})}^{\frac{1}{2}}+T^{-\frac{1}{2}}$
$\lesssim S^{\frac{1}{4}}T^{-\frac{1}{2}}\|T^{-1}\Omega\Gamma^{\leq 8}\psi_{-}\|_{L^{2}(C_{TS})}^{\frac{1}{2}}+T^{-\frac{1}{2}}.$
@@
$\|T^{-1}\Omega\Gamma^{\leq 8}\psi_{-}\|_{L^{2}(C_{TS})}^{\frac{1}{2}}\lesssim\|\Gamma^{\leq 9}\psi_{-}\|_{L^{2}(C_{TS})}^{\frac{1}{2}}\lesssim S^{\frac{1}{4}}.$
@@
$\|\Gamma^{\leq 7}\psi_{-}\|_{L^{2}(C_{TS})}\lesssim ST^{-\frac{1}{2}},$
@@
$\|\Gamma^{\leq 6}\psi_{+}\|_{L^{2}(C_{TS})}\lesssim ST^{-\frac{1}{2}},$
@@
At this stage we are ready to consider separately the interior case $C_{TS}^{+}$ and the exterior case $C_{TS}^{-}$
@@
For this density we first estimate its integral
@@
$\int_{C_{TS}}|\partial_{r}e|\,dxdt\lesssim S.$
Now we think of $C_{TS}^{+}$ as being foliated by hyperboloids, with the transversal direction given by the radial direction. In the radial direction the set $C_{TS}^{+}$ has thickness $O(S)$, so the two estimates above allow us to bound the trace of $e$ on hyperboloids,
@@
The $\psi$ coefficient on the left is larger than the one on the right by a factor of $S$. So reiterating with a slightly wider bump $\chi$ we arrive at
@@
$\|\psi_{+}\|_{L^{2}(H\cap C_{TS}^{-})}^{2}\lesssim\|\psi_{+}\|_{L^{2}(C_{TS}^{-})}\|\partial_{r}\psi_{+}\|_{L^{2}(C_{TS}^{-})}\lesssim\frac{1}{S^{\frac{1}{2}}T^{\frac{1}{2}}}\frac{1}{S}$
@@
$E_{H}(A):=\int_{H\cap C_{T}}(\nu_{0}-|\nu_{x}|)|(\partial_{t}-\partial_{r})A|^{2}+(\nu_{0}+|\nu_{x}|)(|(\partial_{t}+\partial_{r})A|^{2}+|/\!\!\!\!\nabla A|^{2})\,d\sigma,$
@@
This already suffices to get the pointwise bounds for $\mathscr{T}A_{\lambda}$, for which we write
$\partial_{r}\mathscr{T}A_{\lambda}=r^{-1}\partial_{r}\Omega A_{\lambda}.$
Then using twice Lemma 5.7 we obtain
@@
(5.39) $(t^{2}-r^{2})\partial_{r}^{2}A_{\lambda}+\frac{1}{r}\Omega\nabla A_{\lambda}+\frac{1}{r}\nabla A_{\lambda}=G_{\lambda}.$
@@
$\|\Omega^{\leq 2}\partial_{r}A_{\lambda}\|_{L^{\frac{5}{3}}(H\cap C_{TS})}^{\frac{5}{3}}$ $\lesssim\|\Omega^{\leq 2}\partial_{r}A_{\lambda}\|_{L^{2}(C_{TS})}^{\frac{2}{3}}\|\Omega^{\leq 2}\partial_{r}^{2}A_{\lambda}\|_{L^{\frac{3}{2}}(C_{TS})}+S^{-1}(ST^{3})^{\frac{1}{6}}\|\Omega^{\leq 2}\partial_{r}A_{\lambda}\|_{L^{2}(C_{TS})}^{\frac{5}{3}}$

### v5

@@ 4.5. Bounds for a mix of derivatives and vector fields
    (4.24) $\|\Omega\partial^{J}\phi\|_{L^{p}_{t,x}(C_{T}^{+})}\lesssim\|\Omega^{\leq 3}\phi\|_{L^{2}_{t,x}(C_{T}^{+})}^{\frac{1}{3}}\|\partial^{\leq\frac{3}{2}|J|}\phi\|_{L^{q}_{t,x}(C_{T}^{+})}^{\frac{2}{3}}.$
@@
    (4.25) $\|\Omega^{2}\partial^{J}\phi\|_{L^{p}_{t,x}(C_{T}^{+})}\lesssim\|\Omega^{\leq 3}\phi\|_{L^{2}_{t,x}(C_{T}^{+})}^{\frac{2}{3}}\|\partial^{\leq 3|J|}\phi\|_{L^{q}_{t,x}(C_{T}^{+})}^{\frac{1}{3}}.$
@@
(4.27) $T_{1}^{\frac{1}{2}}\|F_{1,I}\|_{L^{2}_{t,x}(C_{T_{1}})}\lesssim\varepsilon\,(\|\Gamma^{\leq 9}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}+\|\Gamma^{\leq 9}A\|_{L^{\infty}_{t}L^{3}_{x}(C_{T_{1}})}),$
(4.28) $T_{1}^{\frac{1}{3}}\|G_{1,I}\|_{L^{1}_{t}L^{\frac{3}{2}}_{x}(C_{T_{1}})}\lesssim\varepsilon\,\|\Gamma^{\leq 9}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}.$
@@
$T_{1}^{\frac{1}{2}}\|\partial^{I-I_{1}}A_{\mu}\gamma^{\mu}\Omega\partial^{I_{1}}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}\lesssim T_{1}^{\frac{1}{2}}\|\partial^{I-I_{1}}A\|_{L^{p_{1}}_{t,x}(C_{T_{1}})}\|\Omega\partial^{I_{1}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}.$
@@
$\|\partial^{I-I_{1}}A\|_{L^{p_{1}}_{t,x}(C_{T_{1}})}$ $\lesssim\|\partial^{\leq\frac{3}{2}|I|}A\|_{L^{3}_{t,x}(C_{T_{1}})}^{\theta_{1}}\|A\|_{L^{\infty}_{t,x}(C_{T_{1}})}^{1-\theta_{1}}$
$\lesssim T_{1}^{\frac{1}{3}\theta_{1}}\|\partial^{\leq 9}A\|_{L^{\infty}_{t}L^{3}_{x}(C_{T_{1}})}^{\theta_{1}}(C_{0}\varepsilon{T_{1}}^{-1})^{1-\theta_{1}},$
@@
$\|\Omega\partial^{I_{1}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}\lesssim\|\Omega^{\leq 3}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}^{\frac{1}{3}}\|\partial^{\leq\frac{3}{2}|I_{1}|}\psi\|_{L^{q_{2}}_{t,x}(C_{T_{1}})}^{\frac{2}{3}}.$
@@
$\|\partial^{\leq\frac{3}{2}|I_{1}|}\psi\|_{L^{q_{2}}_{t,x}(C_{T_{1}})}\lesssim\|\partial^{\leq\frac{3}{2}|I|}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}^{\theta_{2}}\|\psi\|_{L^{6}_{t,x}(C_{T_{1}})}^{1-\theta_{2}}.$
@@
$\|\Omega\partial^{I_{1}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}$ $\lesssim T_{1}^{\frac{1}{6}}\|\Omega^{\leq 3}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}^{\frac{1}{3}}T_{1}^{\frac{1}{2}\theta_{2}\cdot\frac{2}{3}}\|\partial^{\leq 9}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}^{\frac{2}{3}\theta_{2}}$
@@
(4.29) $T_{1}^{\frac{1}{2}}\|\partial^{I-I_{1}}A_{\mu}\gamma^{\mu}\Omega\partial^{I_{1}}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}\lesssim C_{0}\varepsilon(\|\Gamma^{\leq 9}\psi\|_{L^{\infty}_{t}L^{2}_{x}(C_{T_{1}})}+\|\Gamma^{\leq 9}A\|_{L^{\infty}_{t}L^{3}_{x}(C_{T_{1}})}).$
@@
$T_{1}^{\frac{1}{2}}\|\partial^{I-I_{1}}A_{\mu}\gamma^{\mu}\partial_{\nu}\Omega\partial^{I_{2}}\psi\|_{L^{2}_{t,x}(C_{T_{1}})}\lesssim T_{1}^{\frac{1}{2}}\|\partial^{I-I_{1}}A\|_{L^{p_{1}}_{t,x}(C_{T_{1}})}\|\partial_{\nu}\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}$
@@
$\|\partial_{\nu}\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}\lesssim\|\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}^{\frac{1}{2}}\|\partial^{\leq 2}\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}^{\frac{1}{2}}.$
@@
$\|\partial_{\nu}\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}\lesssim\|\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}+\|\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}^{\frac{1}{2}}\|\partial^{2}\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}^{\frac{1}{2}}.$
Using again that the commutator of $\Omega$ with regular derivatives yields regular derivatives which can be dealt with by Gagliardo-Nirenberg estimates, it remains to estimate $\|\Omega\partial^{I_{2}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}$ and $\|\Omega\partial^{2}\partial^{I_{2}}\psi\|_{L^{q_{1}}_{t,x}(C_{T_{1}})}$. But as $|I_{2}|$ is even, we can proceed here as in the case $|I_{1}|$ even. In conclusion, we also obtain (4.29) in the case $|I_{1}|$ odd.
@@ 5. Pointwise bounds
Our main result here is linear (applies to (4.5)) and applies at a fixed dyadic scale $T$. Because of this, we omit the $T^{C\varepsilon}$ factor in Proposition 4.1. Then we want to show that
@@
(5.2) $|\psi|\lesssim\langle t+r\rangle^{-\frac{3}{2}}\langle t-r\rangle_{-}^{-\delta},$
@@
[…]
@@
(5.8) $T^{\frac{1}{2}}\|\Gamma^{\leq 2k}F\|_{L^{2}_{t,x}(C_{T})}+T^{\frac{1}{3}}\|\Gamma^{\leq 2k}G\|_{L^{\frac{3}{2}}_{t,x}(C_{T})}\lesssim 1.$
@@
(5.15) $\|\Gamma^{\leq k}\psi_{+}\|_{L^{2}_{t,x}(C_{TS})}\lesssim$ $\ T^{\frac{1}{2}}\|\Gamma^{\leq k}\psi\|_{X_{T}},$
$\|\Gamma^{\leq k}\psi_{-}\|_{L^{2}_{t,x}(C_{TS})}\lesssim$ $\ S^{\frac{1}{2}}\|\Gamma^{\leq k}\psi\|_{X_{T}}.$
@@
(5.16) $\|P^{\theta}_{+}\Gamma^{\leq k}\psi\|_{L^{2}_{t,x}(C_{TS})}\lesssim T^{\frac{1}{2}}\|{\Gamma^{\leq k}}\psi\|_{X_{T}},$
@@
(5.17) $\|P_{-}^{\theta}\Gamma^{\leq k}\psi\|_{L^{2}_{t,x}(C_{TS})}\lesssim S^{\frac{1}{2}}\|\Gamma^{\leq k}\psi\|_{X_{T}}.$
@@
(5.21) $\|\Gamma^{\leq 6}F_{1}\|_{L^{2}_{t,x}(C_{TS})}\lesssim T^{-\frac{1}{2}}.$
@@
(5.23) $\|\Gamma^{\leq 9}\psi_{-}\|_{L^{2}_{t,x}(C_{TS})}\lesssim S^{\frac{1}{2}},\quad\|\Gamma^{\leq 9}\psi_{+}\|_{L^{2}_{t,x}(C_{TS})}\lesssim T^{\frac{1}{2}},\quad\|\Gamma^{\leq 6}F_{\pm}\|_{L^{2}_{t,x}(C_{TS})}\lesssim T^{-\frac{1}{2}}.$
@@
(5.24) $\|\Gamma^{\leq 6}\psi_{\pm}\|_{L^{2}_{t,x}(C_{TS})}\lesssim ST^{-\frac{1}{2}}.$
A direct application of (5.22) and (5.23) yields the required estimate for $\Gamma^{\leq 6}\psi_{-}$, but only $\|\Gamma^{\leq 6}\psi_{+}\|_{L^{2}_{t,x}(C_{TS})}\lesssim S^{\frac{1}{2}}$. To improve on that bound, we look at the first equation in (5.22) for $\Gamma^{\leq 7}\psi_{-}$ and note that the corresponding right-hand side is given by
@@
$\|r^{-1}\Omega\Gamma^{\leq 7}\psi_{-}\|_{L^{2}_{t,x}(C_{TS})}$ $\lesssim\frac{1}{T}\|\partial\Omega\Gamma^{\leq 6}\psi_{-}\|_{L^{2}_{t,x}(C_{TS})}+T^{-\frac{1}{2}}$
$\lesssim\frac{1}{T}\|\partial^{\leq 2}\Omega\Gamma^{\leq 6}\psi_{-}\|_{L^{2}_{t,x}(C_{TS})}^{\frac{1}{2}}\|\Omega\Gamma^{\leq 6}\psi_{-}\|_{L^{2}_{t,x}(C_{TS})}^{\frac{1}{2}}+T^{-\frac{1}{2}}$
$\lesssim S^{\frac{1}{4}}T^{-\frac{1}{2}}\|T^{-1}\partial^{\leq 2}\Omega\Gamma^{\leq 6}\psi_{-}\|_{L^{2}_{t,x}(C_{TS})}^{\frac{1}{2}}+T^{-\frac{1}{2}}$
$\lesssim S^{\frac{1}{4}}T^{-\frac{1}{2}}\|T^{-1}\Omega\Gamma^{\leq 8}\psi_{-}\|_{L^{2}_{t,x}(C_{TS})}^{\frac{1}{2}}+T^{-\frac{1}{2}}.$
@@
$\|T^{-1}\Omega\Gamma^{\leq 8}\psi_{-}\|_{L^{2}_{t,x}(C_{TS})}^{\frac{1}{2}}\lesssim\|\Gamma^{\leq 9}\psi_{-}\|_{L^{2}_{t,x}(C_{TS})}^{\frac{1}{2}}\lesssim S^{\frac{1}{4}}.$
@@
$\|\Gamma^{\leq 7}\psi_{-}\|_{L^{2}_{t,x}(C_{TS})}\lesssim ST^{-\frac{1}{2}},$
@@
$\|\Gamma^{\leq 6}\psi_{+}\|_{L^{2}_{t,x}(C_{TS})}\lesssim ST^{-\frac{1}{2}},$
@@
At this stage we are ready to consider separately the interior case $C_{TS}^{+}$ and the exterior case $C_{TS}^{-}$.
@@
For this density, employing (5.24), we first estimate its integral
@@
$\int_{C_{TS}}|\partial_{r}e|\,dxdt\lesssim S,$
where we used (5.24) and the bound (5.23) for $F_{\pm}$. Now we think of $C_{TS}^{+}$ as being foliated by hyperboloids, with the transversal direction given by the radial direction. In the radial direction the set $C_{TS}^{+}$ has thickness $O(S)$, so the two estimates above allow us to bound the trace of $e$ on hyperboloids,
@@
As $|\psi|^{2}=|\psi_{+}|^{2}+|\psi_{-}|^{2}$, the coefficients of $\psi_{+}$ and $\psi_{-}$ on the left are larger than the one on the right by at least a factor of $S$. So reiterating with a slightly wider bump $\chi$ we arrive at
@@
$\|\psi_{+}\|_{L^{2}(H\cap C_{TS}^{-})}^{2}\lesssim\|\psi_{+}\|_{L^{2}_{t,x}(C_{TS}^{-})}\|\partial_{r}\psi_{+}\|_{L^{2}_{t,x}(C_{TS}^{-})}\lesssim\frac{1}{S^{\frac{1}{2}}T^{\frac{1}{2}}}\frac{1}{S}$
@@
$E_{H}(A):=\int_{H\cap C_{T}}(\nu_{0}-|\nu_{x}|)|(\partial_{t}-\partial_{r})A|^{2}+(\nu_{0}+|\nu_{x}|)(|(\partial_{t}+\partial_{r})A|^{2}+|/\penalty 50\!\!\!\!\nabla A|^{2})\,d\sigma,$
@@
We first address the pointwise bounds for $\mathscr{T}A_{\lambda}$. We observe that $\partial_{t}+\partial_{r}=\frac{1}{r}\Omega_{0r}-\frac{t-r}{r}\partial_{r}$ with $\Omega_{0r}=\frac{x_{i}}{r}\Omega_{0i}$. Using the pointwise estimate which we show for $\partial_{r}A$ in (5.43) below, we obtain $\|\frac{t-r}{r}\partial_{r}A\|_{L^{\infty}(H\cap C_{TS})}\lesssim ST^{-1}S^{-\frac{1}{2}}T^{-1}\lesssim T^{-\frac{3}{2}}$. For the remaining components $\frac{1}{r}\Omega$ of the tangential derivatives, we use
(5.36) $\partial_{r}(r^{-1}\Omega A_{\lambda})=r^{-1}\partial_{r}\Omega A_{\lambda}-r^{-2}\Omega A_{\lambda}.$
Using twice Lemma 5.7, we obtain
@@
(5.40) $\frac{(t^{2}-r^{2})}{r^{2}}\partial_{r}^{2}A_{\lambda}+\frac{1}{r}\Omega\nabla A_{\lambda}+\frac{1}{r}\nabla A_{\lambda}=G_{\lambda}.$
@@
$\|\Omega^{\leq 2}\partial_{r}A_{\lambda}\|_{L^{\frac{5}{3}}(H\cap C_{TS})}^{\frac{5}{3}}$ $\lesssim\|\Omega^{\leq 2}\partial_{r}A_{\lambda}\|_{L^{2}_{t,x}(C_{TS})}^{\frac{2}{3}}\|\Omega^{\leq 2}\partial_{r}^{2}A_{\lambda}\|_{L^{\frac{3}{2}}_{t,x}(C_{TS})}$
$\qquad+S^{-1}(ST^{3})^{\frac{1}{6}}\|\Omega^{\leq 2}\partial_{r}A_{\lambda}\|_{L^{2}_{t,x}(C_{TS})}^{\frac{5}{3}}$